OpenStudy (anonymous):

What is the slope of the line 8x + 6y = 12?

OpenStudy (muzzack):

6y=-8x+12 y=-.75+2

OpenStudy (anonymous):

so, you have to change this equation in to slope-intercept form 8x + 6y = 12 6y = -8x + 12 y = (-8x + 12) / (6) y = -1.333x + 2 so the slope intercept form is y = mx + b where 'm' is the slope, and 'b' is the y intercept

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yep >.<

OpenStudy (anonymous):

i'm going to post more, help me out please! :)

OpenStudy (anonymous):

Kay :)